On December 31, 2010, Company Parent issued a 3-year, 10% annual coupon bond with a face value of $100,000.

In 2011, Company Parent purchases digital recording machinery for $55,000. The equipment has an estimated useful life of five years and an estimated salvage value of $11,000. The company expects to produce 20,000 units of output using this machinery, with 6,000 units in each of the first two years, 3,000 units in the next two years, and 2,000 units in the fifth year.

Suppose that we are given the following:

  • December 31, 2014 Company Parent (the investor) invests $10,000 in return for 28% of the common shares of Company New (the investee).
  • During 2015, Company New earns $5,000 and pays dividends of $1,000.
  • During 2016, Company New earns $7,000 and pays dividends of $1,500.

Suppose that on January 1, 2016, Company Parent acquires 80% of the common stock of Company Subsidiary by paying $18,000 in cash to the shareholders of Company Subsidiary. The pre-acquisition income statements are as follows:

Income statements

December 31, 2010

Company Parent

Company Subsidiary

Revenue

$70,000

$30,000

Expenses

-40,000

-16,000

Net income

$30,000

$14,000

Dividends paid

$9,000

1.Prepare amortization table for the bond issued on December 31, 2010. Assume that the bond was issued at a market rate of interest of 9%. Please explain your calculations

2.Prepare amortization table for the bond issued on December 31, 2010. Assume that the bond was issued at a market rate of interest of 11%. Please explain your calculations.
Answer #1

Answer 1:

Face value of bond = $100,000

Coupon rate = 10%

Annual coupon bond, coupon amount = $100,000 * 10% = $10,000

Maturity = 3 years

Market rate of interest = 9%

To get issue price we will use PV function of excel:

PV (rate, nper, pmt, fv, type)

= PV (9%, 3, -10000, -100000, 0)

= $102,531.29

Issue price = $102,531.29

As market interest rate is lower than the coupon rate, bond is issued at premium. The premium is to be amortized over life of the bond.

Bond Premium = $102,531.29 - $100,000 = $2,531.29

Year 1 (2011) interest expense = 102531.29 * 9% = $9,227.82

Year 1 amortization of premium = $10,000 - $9227.82 = $772.18

Book value at the end of year 1 = $102,531.29 - $772.18 = $101,759.11

Similarly year 2 and year 3 calculations are done and given below in amortization schedule.

Amortization schedule is as follows:

Interest payment Interest expense (9% * Amortization of Bond (10% * Face value) Previous BV in col. F ) Premium (B-C) Bond Premium $2,531.29 $102,531.29 $1,759.11 $101,759.11 $917.43$100,917.43 Date Book Value 31-Dec-10 31-Dec-11$10,000.00 30-Dec-12 $10,000.00 30-Dec-13 $10,000.00 Total $9,227.82 $9,158.32 $9,082.57 $27,468.71 $772.18 $841.68 $917.43 $2,531.29 $0.00 $100,000.00 $30,000.00

Answer 2:

Face value of bond = $100,000

Coupon rate = 10%

Annual coupon bond, coupon amount = $100,000 * 10% = $10,000

Maturity = 3 years

Market rate of interest = 11%

To get issue price we will use PV function of excel:

PV (rate, nper, pmt, fv, type)

= PV (11%, 3, -10000, -100000, 0)

= $97,556.29

Issue price = $97,556.29

As market interest rate is higher than the coupon rate, bond is issued at discount. The discount is to be amortized over life of the bond.

Bond Discount = $100,000 - $97,556.29 = $2,443.71

Year 1 (2011) interest expense = 197556.29 * 11% = $10,731.19

Year 1 amortization of discount = $10,731.19 - $10,000 = $731.19

Book value at the end of year 1 = $$97,556.29 + $731.19 = $98,287.48

Similarly year 2 and year 3 calculations are done and given below in amortization schedule.

Amortization schedule is as follows:

Interest payment Interest expense (11% Amortization of Bond | Bond (10% * Face value) * Previous BV in col. F ) Premium (CB) Date Book Value Discount 31-Dec-10 31-Dec-11$10,000.00 30-Dec-12 $10,000.00 30-Dec-13$10,000.00 Total $10,731.19 $10,811.62 $10,900.90 $32,443.71 $731.19 $811.62 $900.90 $2,443.71 $2,443.71 $97,556.29 $1,712.52 $98,287.48 $900.90 $99,099.10 $0.00$100,000.00 $30,000.00